Main INDEX
Monthly INDEX
PREV
NEXT
User name Ulmer
Log entry time 00:11:43 on November 6,2002
Entry number 89709
keyword=Swing Shift Summary
Start shift with reasonably stable beam and Q3D_g40 kinematics. Asked to go to 110 uA, but beam stability became worse again. So stay at 100 uA current.
21:30: Finish Q3D_g40 with 205 counts, instead of runplan 280 counts.
21:33: MCC calls just before we begin pointing run, to say our energy feedback was OFF, though position feedback was ON. I thought I was tracking 1C12 X (and it showed flat and near zero line), but in fact I was looking at earlier data. Sorry. In any case, energy feedback is now ON.
21:45: Pointing run for Q3D_g40 finished.
21:50: Set fields for RIGHT arm for Q3D_g50 and did trigsetup.
22:00: Could not move RIGHT spectrometer. Bogie PS did not come up from GUI.
Called Ed Folts. Ed described procedure; we took controlled access to recycle bogie supplies. It didn't work. Ed has come in to rescue us again. Apparently, the control board for the middle bogies is having problems. Ed disconnected and reconnected connectors to the board and that fixed the problem. So we changed angles at 00:00 and are ready to start pointing run for Q3D_g50. At this point, MCC calls and reports problems in injector....